OpenStudy (anonymous):

anyone willing to do a 25 question quiz

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

send me your email then :]

OpenStudy (anonymous):

i dont share emails, maybe you are a STALKER, I knewer know

OpenStudy (anonymous):

lol well i can't do it any other way

OpenStudy (anonymous):

just post the questions here...

OpenStudy (anonymous):

i can't it won't let me i need an email sorry man

OpenStudy (anonymous):

it will let you, just attach the file here

OpenStudy (anonymous):

file with questions

OpenStudy (anonymous):

1. Find the slope of a line. A) B) C) D) 2. Find the slope of a line. A) B) C) D) 3. Find the slope of a line. A) B) undefined C) D) 0 4. Find the slope of the line through the pair of points. C(5, 0), M(–3, –2) A) B) C) D) 5. Find the slope of the line through the pair of points. A(2, –2), D(–2, –5) A) B) C) D) 6. Find the slope of the line through the pair of points. C(–2, 0), P(3, 1) A) B) C) D) 7. Find the slope of the line through the pair of points. F(–4, 5), D(2, –3) A) B) C) D) 8. Identify the slope and y-intercept of the graph of the equation. Then graph the equation. y = x + 3 A) slope: ; y-intercept: –3 B) slope: ; y-intercept: –3 C) slope: ; y-intercept: 3 D) slope: ; y-intercept: 3 9. Identify the slope and y-intercept of the graph of the equation. Then graph the equation. y = –2x A) slope: 2; y-intercept: 0 B) slope: –2; y-intercept: 1 C) slope: undefined; y-intercept: none D) slope: –2; y-intercept: 0 10. Identify the slope and y-intercept of the graph of the equation. Then graph the equation. y = x – 2 A) slope: ; y-intercept: 2 B) slope: ; y-intercept: –2 C) slope: ; y-intercept: 2 D) slope: ; y-intercept: –2 11. The temperature on a particular day started at 8°F. It decreased steadily by 4° each hour. The function y = 8 – 4x models the temperature, where x is the number of hours and y is the temperature. Graph the equation. A) B) C) D) 12. The temperature on a particular day started at 9°F. It decreased steadily by 2° each hour. The function y = 9 – 2x models the temperature, where x is the number of hours and y is the temperature. Graph the equation. A) B) C) D) 13. The temperature on a particular day started at 24°F. It decreased steadily by 4° each hour. The function y = 24 – 4x models the temperature, where x is the number of hours and y is the temperature. Graph the equation. A) B) C) D) 14. The temperature on a particular day started at 7°F. It rose steadily by 2° each hour. The function y = 7 + 2x models the temperature, where x is the number of hours and y is the temperature. Graph the equation. A) B) C) D) 15. Sam has a home-based business selling custom-designed jewelry at home parties. He charges $25 to design the party and then $8 per guest. Write a function rule that relates the total cost of the party to the number of guests, (n). A) f(n) = 8 + 25n B) f(n) = 8n – 55 C) f(n) = 25 + 8n D) f(n) = 8 – 55n 16. Mike has a home-based business hosting corporate meet-and-greet parties. He charges $30 to design the party and then $9.5 per guest. Write a function rule that relates the total cost of the party to the number of attendees, (n). A) f(n) = 9.5 – 55n B) f(n) = 9.5 + 30n C) f(n) = 9.5n – 55 D) f(n) = 30 + 9.5n 17. Nora has a home-based business selling custom-designed jewelry at home parties. She charges $50 to design the party and then $8 per guest. Write a function rule that relates the total cost of the party to the number of guests, (n). A) f(n) = 8 + 50n B) f(n) = 50 + 8n C) f(n) = 8n – 55 D) f(n) = 8 – 55n 18. The sales tax in one state is 4%. Write a function rule for finding the total cost of an item with selling price x. Then find the total cost of a CD player with a sales price of $200. A) f(x) = 200x – 4; $198.96 B) f(x) = 0.04x + 200; $208 C) f(x) =1.04x; $8 D) f(x) = 4 + 200x; $201.04 19. The sales tax in one state is 4%. Write a function rule for finding the total cost of an item with selling price x. Then find the total cost of a pair of the latest endorsed athletic shoes with a sales price of $230. A) f(x) = 4 + 230x; $231.04 B) f(x) =1.04x; $9.2 C) f(x) = 230x – 4; $228.96 D) f(x) = 0.04x + 230; $239.2 20. The sales tax in one state is 6%. Write a function rule for finding the total cost of an item with selling price x. Then find the total cost of a CD player with a sales price of $270. A) f(x) = 6 + 270x; $271.06 B) f(x) =1.06x; $16.2 C) f(x) = 270x – 6; $268.94 D) f(x) = 0.06x + 270; $286.2 21. The sales tax in one state is 7%. Write a function rule for finding the total cost of an item with selling price x. Then find the total cost of a pair of the latest endorsed athletic shoes with a sales price of $250. A) f(x) = 0.07x + 250; $267.5 B) f(x) = 7 + 250x; $251.07 C) f(x) = 250x – 7; $248.93 D) f(x) =1.07x; $17.5 22. Write a rule for the linear function in the graph. A) y = x – 1 B) y = x + 1 C) y = x + 1 D) y = x + 1 23. Write a rule for the linear function in the graph. A) y = x + 1 B) y = x – 1 C) y = x – 1 D) y = x – 1 24. Write a rule for the linear function in the graph. A) y = x – 1 B) y = x – 1 C) y = x + 1 D) y = x – 1 25. Write a rule for the linear function in the graph. A) y = x + 15 B) y = x + 15 C) y = x + 15 D) y = x – 15

OpenStudy (anonymous):

good luck! please good answers and try to do effecint work in a good amount of time!